OpenStudy (peachytea):

the average of six numbers is 22. the numbers are 20, 17, 24, 18, 29, and x. find x

OpenStudy (anonymous):

(20+17+24+18+29+x)÷6=22

OpenStudy (anonymous):

solve for x

OpenStudy (jdoe0001):

\(\large { \begin{array}{rllll} 20\\17\\243\\18\\29\\+x \\\hline\\ \square ? \end{array}\qquad \cfrac{\square }{6}=22\impliedby average \\ \quad \\ \cfrac{20+17+243+18+29+x}{6}=22\impliedby \textit{solve for "x"} }\)

OpenStudy (anonymous):

to cancel out the 6, multiply both sides by 6

OpenStudy (anonymous):

then simplify both sides of the equation

OpenStudy (anonymous):

to get x+108=132

OpenStudy (anonymous):

I think you can do the rest by yourself :)

OpenStudy (peachytea):

24!

OpenStudy (anonymous):

yes